Within a computer service company, 20 percent of the male employees and 30 percent of the female employees attend night school. If 40 percent of all employees are male what percent of all the employees attend night school? 20% 030% O25% 27.596 26% Clear The average length of time that three lawn mowers operate on a given day is 930 minutes. None of the mowers operates for less than 900 minutes. What is the greatest possible length of time any of the mowers operates? 0950?990 0960 0930 0900 Clear In a work-out facility, there are two racks with 40 towels per rack. On Tuesday, there were 30 towels in use, on Wednesday s0, and on Thursday 25. what fraction, on average is in use for the 3 days? Clear 8 A company rented a copier for $600 per month. Five years later, an accountant calcuiated that f the company had purchased the copier and pald a s100 monthly maintenance charge, the company would have saved $700. What was the purchase price of the copier? Ciear o$27,250 0$25,00& 0 $30,000 0$29,300 O$26,000 Next Page
Answer #1

Ans A: 26%.

Since 40% are male, 40 out of 100 employees are male and rest i.e. 60 employees are female.
Of the 40, 8 males go to night school and 30% of females i.e. 18 females go to night school.
So, a total 26 employees out of 100 employees i.e. 26% of employees attend night school.

Ans B. 990 minutes

Three lawn mowers operate for 930 minutes per day 900.
i.e. the three operate for a total of 3 x 930 = 2790 minutes.
Out of these if at least two of them mow for a minimum 900 minutes i.e.1800 minutes in total,
the third would be mowing for 2790 - 1800 = 990 minutes.

Ans C. 7/16

Total towels used in the thre days = 30 + 50 + 25 = 105.
So on an average 105 / 3 = 35 towels are in use.
i.e. 35 out of 80 i.e. 35 / 80 = 7/16 are in use during the 3 days.

Ans D. $29300

Total rent paid on the copier for 5 years i.e. 60 months
= 60 months x $600 = $36000.
Had it been purchased there would have been a maintenance charge of 60 x 100 = $6000 for the 5 years.
Also there would been a saving of $700.
So the copier would have cost = $36000 - $6000 - $700 = $29300.